# Ali Dasdan KD Consulting Saratoga, CA, USA

### v1

@@ Abstract
In 1930, John Maynard Keynes predicted that technological progress would reduce the working week to 15 hours by century’s end. Despite productivity increasing eightfold since then, workers globally still work roughly double these hours. Separately, leading AI researchers warn of existential risks from artificial intelligence, yet development accelerates rather than slows. We demonstrate these failures share identical game-theoretic structure: coordination failures where individually rational choices produce collectively suboptimal outcomes.
We synthesize from established game theory five necessary and sufficient conditions that characterize such coordination failures as structural tragedies: N-player structure with multiple independent actors, binary choices generating negative externalities, dominance property where defection yields higher payoffs, Pareto-inefficiency where cooperation dominates mutual defection, and enforcement difficulty from structural barriers. We extend this framework to quantify tragedy severity through condition intensities, introducing a Tragedy Index that reveals AI governance faces orders-of-magnitude greater coordination difficulty than historical cases including climate change or nuclear weapons. We validate this framework across canonical cases including commons tragedies, security dilemmas, bank runs, and space debris.
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We demonstrate this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence demonstrates that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, productivity welfare decoupling persists, manifesting differently across institutional contexts.
Applied to AI governance, we show development faces the same structure but with amplified intensity across eight dimensions compared to successful arms control cases, making coordination structurally more difficult than for nuclear weapons. The Russia-Ukraine drone war provides empirical validation: despite international dialogue on weapons governance, both sides escalated from zero to thousands of drones monthly within two years, driven by competitive necessity that overwhelmed coordination aspirations.
@@ 1.4 Preview of Contributions
This paper makes three contributions:
@@
Second, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence demonstrates that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
Third, we demonstrate that AI development exhibits the same structure but with amplified intensity. Systematic comparison with historical arms control reveals that AI faces unfavorable conditions on every dimension where nuclear, biological, chemical (NBC) weapons coordination succeeded: dual-use inseparability, verification impossibility, low entry barriers, absence of substitutes, and convergence of economic and military imperatives. The Russia-Ukraine drone war provides empirical validation, showing how rapidly competitive dynamics override governance intentions when existential stakes are involved.
@@ 1.5 Organization
The remainder of this paper is organized as follows. Section 2 reviews the related literature on social dilemmas, productivity, and AI safety. Section 3 introduces our unified framework, defining the five necessary conditions for structural tragedy and the Tragedy Index. Section 4 applies this framework to labor markets, providing a formal proof of the coordination failure in working hours. Section 5 extends the analysis to AI governance, offering a comparative analysis with historical arms control and examining the Russia-Ukraine drone conflict. Section 6 discusses testable predictions, falsifiability, and potential interventions. Section 7 concludes.
@@ 2.1 The Prisoner’s Dilemma and Social Dilemmas
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as $N$ increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
@@
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
The folk theorem literature (Fudenberg and Maskin, 1986) demonstrates that repeated interaction can sustain cooperation under certain conditions through trigger strategies. However, these results require strong assumptions: players must be sufficiently patient, the game must repeat infinitely, and actions must be perfectly observable. When these conditions fail—as we show they do in productivity competition and AI development—cooperation becomes unsustainable even with repetition.
@@ 2.2 Tragedies in Specific Domains
Olson (1965) analyzed collective action problems in public goods provision, showing why rational individuals undercontribute to group benefits. Diamond and Dybvig (1983) modeled bank runs as coordination failures where individual rationality leads to systemic crisis. These analyses demonstrate that coordination failures pervade economic life, not merely environmental or security domains.
@@ 3.1 The Five Conditions
Multiple independent decision-makers ($n\geq 2$) each control their own choices. No single actor can solve the problem unilaterally. Choices are made in a decentralized architecture without centralized coordination.
This condition specifies the basic game structure. If a single actor controlled all decisions, coordination would be trivial. The tragedy emerges from decentralized choice architecture where multiple actors must independently decide their strategies.
@@
Each actor $i$ faces a binary choice between two strategies: Cooperate ($C$) representing restraint, or Defect ($D$) representing advantage-seeking. Individual choices create externalities affecting all actors. Specifically, one actor’s defection imposes negative externalities on others.
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: restrain (cooperate) or expand (defect). The key is that individual choices impose externalities—one actor’s defection makes cooperation more costly for others.
@@
Defection yields higher payoff than cooperation regardless of other actors’ choices. Formally,
@@
Dominance creates the dilemma. Each actor has strict incentive to defect regardless of what others do. This is often driven by existential stakes. In high-stakes tragedies, survival instincts override social norms; if cooperating while others defect leads to elimination (e.g., death, bankruptcy), defection becomes an absolute imperative. When survival is at stake, the incentive to defect becomes overwhelming.
@@
The equilibrium outcome is recognizably suboptimal—all actors would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Actors understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden—it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
@@
Structural barriers prevent effective coordination mechanisms. These barriers may include: (i) anarchy as in no central authority with enforcement power, (ii) verification impossibility as in actions cannot be reliably monitored, (iii) monitoring infeasibility as in costs of monitoring exceed benefits of coordination, or (iv) commitment problems as in actors cannot credibly commit to future cooperation.
@@ 3.2 The Formal Result
Let $s=(s_{1},\ldots,s_{n})$ be a strategy profile where $s_{i}\in\{C,D\}$ for each actor $i$.
First, we show $s^{D}=(D,\ldots,D)$ is a Nash equilibrium. For any actor $i$, their payoff from defecting is $U_{i}(D,s^{D}_{-i})$, where $s^{D}_{-i}$ is the profile where all other players also defect. If actor $i$ unilaterally deviates to cooperation, their payoff becomes $U_{i}(C,s^{D}_{-i})$. By the dominance property (Condition 3), $U_{i}(D,s^{D}_{-i})>U_{i}(C,s^{D}_{-i})$. Therefore, no actor has an incentive to deviate, and $s^{D}$ is a Nash equilibrium.
Second, we show no other profile is a Nash equilibrium. Consider any profile $s$ where at least one actor $j$ cooperates. This actor’s payoff is $U_{j}(C,s_{-j})$, where $s_{-j}$ is the profile of all other players. If actor $j$ deviates to defection, their payoff becomes $U_{j}(D,s_{-j})$. By the dominance property (Condition 3), $U_{j}(D,s_{-j})>U_{j}(C,s_{-j})$. Therefore, actor $j$ has a strict incentive to deviate. Hence, no profile with any cooperation is a Nash equilibrium.
@@
Stability follows immediately from Claim (i). At the equilibrium $s^{D}$, no actor has unilateral incentive to deviate by the dominance property. The equilibrium is thus stable against unilateral deviations.
@@
By Condition 4 (Pareto-inefficiency), $U_{i}(s^{C})>U_{i}(s^{D})$ for all actors $i$. Therefore, universal cooperation strictly Pareto-dominates the Nash equilibrium.
@@
Even though all actors recognize that $s^{C}$ Pareto-dominates $s^{D}$, they cannot coordinate to achieve it. By Condition 5, structural barriers prevent enforcement:
@@
  • If verification is impossible, actors cannot monitor compliance
@@
Without enforcement, any agreement to cooperate is unstable, as each actor has a strict incentive to defect (from Claim (i)). The folk theorem, which allows for cooperation in repeated games, does not apply because its required conditions (perfect monitoring, sufficient patience, infinite repetition) fail.
@@
  • Actors can make credible commitments to future cooperation
Under these conditions, actors can contract to implement $s^{C}$ and sustain it through enforcement mechanisms. Defection can be detected and punished. Commitments become credible. The folk theorem for repeated games may apply: if actors are sufficiently patient and monitoring is adequate, cooperation can be sustained as a subgame perfect equilibrium through trigger strategies that punish defection.
@@ 3.3 Validation Across Canonical Cases
Result: Perfect verification across all canonical cases. Each tragedy satisfies all five conditions. This suggests the framework captures the essential structure that creates coordination failures.
@@ 3.4 Implications of the Framework
Identification: We can identify new tragedies by checking whether a situation satisfies all five conditions. If it does, coordination will be structurally difficult. If any condition fails, standard solutions may suffice.
@@ 3.5 Extension: Condition Intensity and Tragedy Severity
While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. The severity of a structural tragedy increases with the intensity of each condition. This intensity dimension provides additional predictive and diagnostic power.
C1: N-Player Structure - Number and Heterogeneity of Actors.
@@
C2: Externalities - Magnitude and Scope.
@@
C3: Dominance - Strength of Existential Stakes.
@@
  • Medium: Productivity competition—firms face market share loss, not immediate bankruptcy
@@
  • Extreme: AI development with AGI—actors face permanent strategic obsolescence or existential risk
C4: Pareto-Inefficiency - Magnitude of Collective Loss.
@@
  • Medium: Productivity competition—substantial leisure forgone, wage stagnation
  • High: Climate change—potential civilizational collapse
  • Extreme: AI misalignment—potential human extinction
C5: Enforcement Difficulty - Severity of Structural Barriers.
@@ 3.6 Implications of Intensity Analysis
Comparative prediction: We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance faces high intensity on all five conditions simultaneously, while nuclear non-proliferation faced high intensity on only some conditions (dominance, Pareto-inefficiency) but lower intensity on others (entry barriers created enforcement possibility).
@@
Temporal variation explanation: Condition intensities change over time. Productivity competition intensified after 1980 as globalization increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility). This explains why coordination that worked 1945-1980 eroded subsequently.

### v2

@@ Abstract
In 1930, John Maynard Keynes predicted that technological progress would reduce the working week to 15 hours by century’s end. Despite productivity increasing eightfold since then, workers globally still work roughly double these hours. Separately, leading AI researchers warn of existential risks from artificial intelligence, yet development accelerates rather than slows. We demonstrate these failures share identical game-theoretic structure: coordination failures arising when individual rationality produces collective harm.
We synthesize from established game theory five necessary and sufficient conditions that characterize such coordination failures as structural tragedies: N-player structure with multiple independent actors, binary choices generating negative externalities, dominance property where defection yields higher payoffs, Pareto-inefficiency where cooperation dominates mutual defection, and enforcement difficulty from structural barriers. We extend this framework to quantify tragedy severity through condition intensities, introducing a Tragedy Index that reveals governance of transformative AI breakthroughs faces orders-of-magnitude greater coordination difficulty than historical cases including climate change, nuclear weapons, or bank runs. We validate this framework across canonical cases including commons tragedies, security dilemmas, bank runs, and space debris.
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We show this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence confirms that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, productivity welfare decoupling persists, manifesting differently across institutional contexts.
Applied to AI governance, we show development faces the same structure but with amplified intensity across eight dimensions compared to successful arms control cases, making coordination structurally more difficult than for nuclear weapons. The Russia-Ukraine drone war provides empirical validation: despite international dialogue on weapons governance, both sides escalated from dozens to thousands of drones monthly within two years, driven by competitive necessity that overwhelmed coordination aspirations.
@@ 1.4 Preview of Contributions
This paper makes four contributions:
@@
Second, we extend this framework to quantify tragedy severity through condition intensities. While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. We introduce a Tragedy Index that systematically compares coordination difficulty across domains, revealing that governance of transformative AI breakthroughs faces orders-of-magnitude greater challenges than historical cases including climate change, nuclear weapons, or bank runs.
Third, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence shows that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
Fourth, we demonstrate that AI development exhibits the same structure but with amplified intensity. Systematic comparison with historical arms control reveals that AI faces unfavorable conditions on every dimension where nuclear, biological, chemical (NBC) weapons coordination succeeded: dual-use inseparability, verification impossibility, low entry barriers, absence of substitutes, and convergence of economic and military imperatives. The Russia-Ukraine drone war provides empirical validation, showing how rapidly competitive dynamics override governance intentions when existential stakes are involved.
@@ 1.5 Organization
The remainder of this paper is organized as follows. Section 2 reviews the related literature on social dilemmas, productivity, and AI safety. Section 3 introduces our unified framework, defining the five necessary conditions for structural tragedy, proving these conditions are necessary and sufficient, and introducing the Tragedy Index. Section 4 applies this framework to labor markets, providing a formal proof of the coordination failure in working hours and examining the European case as empirical validation. Section 5 extends the analysis to AI governance, offering a comparative analysis with historical arms control and examining the Russia-Ukraine drone conflict. Section 6 discusses testable predictions, falsifiability, and potential interventions. Section 7 concludes.
@@ 2.1 The Prisoner’s Dilemma and Social Dilemmas
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors [^1] We use ’player’ and ’actor’ interchangeably for game participants. face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as $N$ increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
@@
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” (originally coined by Herz, 1950) where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
The folk theorem literature (Fudenberg and Maskin, 1986) proves that repeated interaction can sustain cooperation under certain conditions through trigger strategies. However, these results require strong assumptions: players must be sufficiently patient, the game must repeat infinitely, and actions must be perfectly observable. When these conditions fail—as we show they do in productivity competition and AI development—cooperation becomes unsustainable even with repetition.
@@ 2.2 Tragedies in Specific Domains
Olson (1965) analyzed collective action problems in public goods provision, showing why rational individuals undercontribute to group benefits. Diamond and Dybvig (1983) modeled bank runs as coordination failures where individual rationality leads to systemic crisis. These analyses confirm that coordination failures pervade economic life, not merely environmental or security domains.
@@ 3 A Unified Framework for Structural Tragedies
We use the term “Tragedy” in the strict sense defined by Whitehead (1925) and popularized by Hardin (1968): “the solemnity of the remorseless working of things.” A situation becomes a tragedy not through error or malice, but through inevitability. Specifically, it occurs when individual rationality compels defection (C3) even though actors collectively recognize cooperation as superior (C4), while structural barriers prevent coordination (C5). The tragedy lies in this structural trap: actors are not making mistakes—they are imprisoned by the logic of their own survival.
@@ 3.1 The Five Conditions
Multiple independent decision-makers ($n\geq 2$) each control their own choices. No single player can solve the problem unilaterally. Choices are made in a decentralized architecture without centralized coordination.
This condition specifies the basic game structure. If a single player controlled all decisions, coordination would be trivial. The tragedy emerges from decentralized choice architecture where multiple players must independently decide their strategies.
@@
Each player $i$ faces a binary choice between two strategies: Cooperate ($C$) representing restraint, or Defect ($D$) representing advantage-seeking. Individual choices create externalities affecting all players. Specifically, one player’s defection imposes negative externalities on others.
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: restrain (cooperate) or expand (defect). The key is that individual choices impose externalities—one player’s defection makes cooperation more costly for others.
@@
Defection yields higher payoff than cooperation regardless of other players’ choices. Formally,
@@
Dominance creates the dilemma. Each player has strict incentive to defect regardless of what others do. This is often driven by existential stakes. In high-stakes tragedies, survival instincts override social norms; if cooperating while others defect leads to elimination (e.g., death, bankruptcy), defection becomes an absolute imperative. When survival is at stake, the incentive to defect becomes overwhelming.
@@
The equilibrium outcome is recognizably suboptimal—all players would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Players understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden—it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
@@
Structural barriers prevent effective coordination mechanisms. These barriers may include: (i) anarchy as in no central authority with enforcement power, (ii) verification impossibility as in actions cannot be reliably monitored, (iii) monitoring infeasibility as in costs of monitoring exceed benefits of coordination, or (iv) commitment problems as in players cannot credibly commit to future cooperation.
@@ 3.2 The Formal Result
Let $s=(s_{1},\ldots,s_{n})$ be a strategy profile where $s_{i}\in\{C,D\}$ for each player $i$.
First, we show $s^{D}=(D,\ldots,D)$ is a Nash equilibrium. For any player $i$, their payoff from defecting is $U_{i}(D,s^{D}_{-i})$, where $s^{D}_{-i}$ is the profile where all other players also defect. If player $i$ unilaterally deviates to cooperation, their payoff becomes $U_{i}(C,s^{D}_{-i})$. By the dominance property (Condition 3), $U_{i}(D,s^{D}_{-i})>U_{i}(C,s^{D}_{-i})$. Therefore, no player has an incentive to deviate, and $s^{D}$ is a Nash equilibrium.
Second, we show no other profile is a Nash equilibrium. Consider any profile $s$ where at least one player $j$ cooperates. This player’s payoff is $U_{j}(C,s_{-j})$, where $s_{-j}$ is the profile of all other players. If player $j$ deviates to defection, their payoff becomes $U_{j}(D,s_{-j})$. By the dominance property (Condition 3), $U_{j}(D,s_{-j})>U_{j}(C,s_{-j})$. Therefore, player $j$ has a strict incentive to deviate. Hence, no profile with any cooperation is a Nash equilibrium.
@@
Stability follows immediately from Claim (i). At the equilibrium $s^{D}$, no player has unilateral incentive to deviate by the dominance property. The equilibrium is thus stable against unilateral deviations.
@@
By Condition 4 (Pareto-inefficiency), $U_{i}(s^{C})>U_{i}(s^{D})$ for all players $i$. Therefore, universal cooperation strictly Pareto-dominates the Nash equilibrium.
@@
Even though all players recognize that $s^{C}$ Pareto-dominates $s^{D}$, they cannot coordinate to achieve it. By Condition 5, structural barriers prevent enforcement:
@@
  • If verification is impossible, players cannot monitor compliance
@@
Without enforcement, any agreement to cooperate is unstable, as each player has a strict incentive to defect (from Claim (i)). The folk theorem, which allows for cooperation in repeated games, does not apply because its required conditions (perfect monitoring, sufficient patience, infinite repetition) fail.
@@
  • Players can make credible commitments to future cooperation
Under these conditions, players can contract to implement $s^{C}$ and sustain it through enforcement mechanisms. Defection can be detected and punished. Commitments become credible. The folk theorem for repeated games may apply: if players are sufficiently patient and monitoring is adequate, cooperation can be sustained as a subgame perfect equilibrium through trigger strategies that punish defection.
@@ 3.3 Validation Across Canonical Cases
Result: Perfect verification across all canonical cases. Each tragedy satisfies the complete set of conditions. This suggests the framework captures the essential structure that creates coordination failures.
@@ 3.4 Implications of the Framework
Identification: We can identify new tragedies by checking whether a situation satisfies the framework. If it does, coordination will be structurally difficult. If any condition fails, standard solutions may suffice.
@@ 3.5 Extension: Condition Intensity and Tragedy Severity
While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. The severity of a structural tragedy increases with the intensity of each condition. This intensity dimension provides additional predictive and diagnostic power by distinguishing between manageable nuisances and existential threats.
3.5.1 C1: N-Player Structure - Number and Heterogeneity of Actors.
@@
  • Extreme: Antibiotics resistance with billions of patients, millions of prescribers across medical and agricultural sectors globally—coordination structurally impossible at individual decision level. Bank runs with millions of independent depositors making simultaneous withdrawal decisions during panic.
3.5.2 C2: Externalities - Magnitude and Scope.
@@
  • Extreme: Security dilemmas—states respond to perceived threats before they materialize, creating anticipatory arms races. Intelligence about potential enemy programs triggers immediate counter-programs based on worst-case assumptions, with no waiting period for actual capability development.
3.5.3 C3: Dominance - Strength of Existential Stakes.
@@
  • Medium: Productivity competition—firms face market share loss and competitive pressure, but not immediate bankruptcy. European firms reduced hours and survived, demonstrating stakes are substantial but not existential. Current AI competition exhibits similar dynamics: Google recovered from Transformer lag, multiple competitors remain viable despite capability gaps (ChatGPT, Gemini, DeepSeek).
@@
  • Extreme: Transformative AI (AGI/superintelligence)— actors face permanent strategic obsolescence with no recovery possibility, or existential risk from misaligned systems. Unlike current competition where multiple players survive capability gaps, transformative breakthroughs may create winner-take-all dynamics where falling behind means permanent irrelevance.
3.5.4 C4: Pareto-Inefficiency - Magnitude of Collective Loss.
@@
  • Medium: Productivity competition—substantial leisure forgone, wage stagnation. Significant quality-of-life loss but society continues functioning with material prosperity.
  • High: Regional conflicts—major destruction but potential for recovery
  • Extreme: Climate change—tipping points create irreversible changes over millennia; ecosystem collapse and mass extinction cannot be undone on human timescales. Nuclear security dilemmas—potential human extinction from nuclear war. AGI misalignment—potential permanent loss of human control or extinction.
3.5.5 C5: Enforcement Difficulty - Severity of Structural Barriers.
@@ 3.6 Implications of Intensity Analysis
Comparative prediction: We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance (for transformative breakthroughs) faces extreme intensity on three conditions simultaneously (C3, C4, C5), while severe historical cases like nuclear security dilemmas faced extreme intensity on two conditions (C2, C4) and climate change on one (C4). Current AI capabilities competition exhibits high but not extreme intensity, as evidenced by multiple viable competitors despite capability gaps.
@@
Temporal variation explanation: Condition intensities change over time. Productivity competition shows moderate base intensity (C2, C3 at medium due to gradual market dynamics and non-existential stakes), but globalization after 1980 increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility), making even this moderate tragedy harder to address. Climate change intensified as C4 shifted from high to extreme with growing evidence of irreversible tipping points.